The volume of a certain bacterial cell is 2.95 μm3. Give your answers in scientific notation. (a) What is its volume in cubic millimeters (mm3)? (b) What is the volume of (exactly) 1 × 106 cells in liters (L)?
Answer #1

a) We know, 1m = 10^-6 m

then, 1m^3 = 10^-18 m^3

so, 2.95m^3 = 2.95*10^-18 m^3

Whereas, 1 m = 10^3 mm

1m^3 = 10^9 mm

then, 2.95m^3 = 2.95*10^-18 m^3 = 2.95*10^-9 mm^3

answer for a) is 2.95*10^-9 mm^3.   

b) Given, 1 cell volume = 2.95m^3

1m^3 = 10^-18 m^3

then, 1 cell volume = 2.95m^3 = 2.95*10^-18 m^3  

we know, 1m^3 = 1000 L then, volume of 1 cell = 2.95*10^-15 L

So, volume of 1*10^6 cells = 2.95*10^-15*10^6 L = 2.95*10^-9 L

The volume of 1*10^6 cells is 2.95*10^-9 L.